OpenStudy (anonymous):

What is the line that passes through the points (-10, -4) and (-1, 2)

OpenStudy (uri):

You find slope. DO you know how to? :) @shymc7068

OpenStudy (anonymous):

\[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (uri):

@iambatman ._.

OpenStudy (anonymous):

Ok so you need to find the slop first! then you have to use the point slop form. ( Y - Y ) = slope ( X - X ) like the slope is 6/11. equation is ( Y - (-10) ) = 6/11 ( X - (-4) ) than you can find this answer!

OpenStudy (anonymous):

M is your slope, and once you get that you can use the point - slope form if you want and plug in the first coordinate, and form it into y = mx+b form using some algebra. \[y-y _{1}=m(x-x _{1})\]